OpenStudy (anonymous):

solve: 4(x + 1) = 4x + 4

OpenStudy (anonymous):

take 4 to that isde

OpenStudy (anonymous):

as x terms will cancel on both sides x can be any val

OpenStudy (anonymous):

so....x+1 = x +1 = 0

OpenStudy (anonymous):

4x+4 = 4x+4 you can subtract 4 from both sides and get 4x=4x but that is as far as you can go to solve for x. so in reality x can be any number.